OpenStudy (migitmack):

Explain how you would find the solution of 3/sqrt(8)

OpenStudy (anonymous):

3/64 is answer

OpenStudy (migitmack):

but EXPLAIN IS THE KEY WORD

OpenStudy (migitmack):

@Lifegood12 plz explain how u got your answer

OpenStudy (anonymous):

It's explainition is as we has Squrt8 Mean's Square of 8 And 3 is numerater

OpenStudy (migitmack):

someone else has a better answer?

OpenStudy (anonymous):

Sure I'll help you out @Migitmack , although I'm not sure what you're stuck on, I'll show some intermediary steps: As an expression to simplify: \[\frac{3}{\sqrt{8}} = \frac{3}{\sqrt{2*2*2}}=\frac{3}{\sqrt{2*2}*\sqrt{2}}=\frac{3}{2\sqrt{2}}\approx 1.0606601717798212866\] @Lifegood12 Read this before you get yourself in trouble: http://openstudy.com/code-of-conduct In short, medals for good answers & questions, and don't only post the just answer, if you can help the person understand what to do that's better. Also you're wrong. It's not 6/8\(^2\), it was the square-root of 8.

OpenStudy (migitmack):

WOW THX @agentx5 that what i meant by explain

OpenStudy (anonymous):

np, I can understand when you're new to a place you don't know the rules very well, but hopefully I've spread that knowledge onto two new people :-)